The diagonal of a rectangle measures 10. If the rectangle is twice as long as it is wide, then find its dimensions
Answer by checkley77(12844)

x^2+(2x)^2=10
x^2+4x^2=10
5x^2=10
x^2=10/5
x^2=2
x=sqrt2
x=1.414 ans. for the shortest side.
2*1.414=2.828 ans. for the longer side.